Two resistors of 3ohm and 6ohm are connected in parallel. This combination in turn is connected in series with a  2ohm resistor and a 6volt battery. The current in the 2ohm resistor is

Grade:12th pass

2 Answers

Arun
4 years ago
Re=(6×3/9)+2=4 ohms
V= 6 Volts
So,I= V/R
=6/4
= 1.5 A
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
Khimraj
4 years ago

equivalent resistance = 2+2 = 4
I= V/R
=6/4
= 1.5 A

….....................................................................................